Mr. John Backster, a retired executive, desires to invest a portion of his assets in rental property. He has narrowed his choices to two apartment complexes, Windy Acres and Hillcrest Apartments. The anticipated annual cash inflows from each are as follows:

Windy
Acres
Hillcrest
Apartments
  Yearly Aftertax
Cash Inflow
Probability   Yearly Aftertax
Cash Inflow
Probability
30,000 0.2 35,000 0.4
35,000 0.2 40,000 0.2
50,000 0.2 50,000 0.1
65,000 0.2 60,000 0.3
70,000 0.2

Mr. Backster is likely to hold the apartment complex of his choice for about 25 years and will use this period for decision-making purposes. Either apartment can be purchased for $200,000. Mr. Backster uses a risk-adjusted discount rate approach when evaluating investments. His scale is related to the coefficient of variation (for other types of investments, he also considers other measures).

Coefficient of Variation Discount Rate
0–0.35     5%
0.35–0.40     9    (cost of capital)
0.40–0.50     14   
Over 0.50     not considered

a. Compute the risk-adjusted net present value for Windy Acres and Hillcrest Apartments. (Round "PV Factor" to 3 decimal places. Enter the answers in thousands of dollars. Do not round intermediate calculations. Round the final answers to nearest whole dollar.)

Net present value
  Windy Acres $   
  Hillcrest Apartments $   

b-1. Which investment should Mr. Backster accept if the two investments are mutually exclusive?

  • Hillcrest

  • Windy Acres

  • Both

  • None

b-2. Which investment should Mr. Backster accept If the investments are not mutually exclusive and no capital rationing is involved?

  • Windy Acres

  • Hillcrest Apartments

  • Both

  • None
Answer #1
Windy Acres Hilcrest Apartments year PV Factor @ 5% (1/(1.05^n))
Cashflow (x) Probability (p) xp p*((x-mean)^2) Cashflow (x) Probability (p) xp p*((x-mean)^2) 1 0.952
30000 0.2 6000 80000000 35000 0.4 14000 40000000 2 0.907
35000 0.2 7000 45000000 40000 0.2 8000 5000000 3 0.864
50000 0.2 10000 0 50000 0.1 5000 2500000 4 0.823
65000 0.2 13000 45000000 60000 0.3 18000 67500000 5 0.784
70000 0.2 14000 80000000 6 0.746
Mean= sum of x*p 50000 250000000 Mean= sum of x*p 45000 115000000 7 0.711
8 0.677
Std. Deviation Std. Deviation 9 0.645
=(sum of p*((x-mean)^2))^1/2 =(sum of p*((x-mean)^2))^1/2 10 0.614
=250000000^(1/2) =115000000^(1/2) 11 0.585
15811.39 10723.81 12 0.557
13 0.530
Co-efficient of variation Co-efficient of variation 14 0.505
=Std. Deviation/Mean =Std. Deviation/Mean 15 0.481
=15811.39/50000 =10723.81/45000 16 0.458
0.32 0.24 17 0.436
18 0.416
Thus, Discount Rate= Thus, Discount Rate= 19 0.396
5% 5% 20 0.377
21 0.359
Risk Adjusted NPV Risk Adjusted NPV 22 0.342
Year Expected Cashflow PV Factor @ 5% PV Year Expected Cashflow PV Factor @ 5% PV 23 0.326
0 -$2,00,000 1 -$2,00,000.000 0 -$2,00,000 1 -$2,00,000.000 24 0.310
1-25 $50,000 14.094 $7,04,697.228 1-25 $45,000 14.094 $6,34,227.505 25 0.295
NPV $5,04,697 NPV $4,34,228 14.094

b-1 if the two investments are mutually exclusive, Mr. Backster should accept Windy Acres as it has higher NPV.

b-2 If the investments are not mutually exclusive and no capital rationing is involved, Mr. Backster should accept Both as both have Positive NPV.

(Kindly Rate Positively, if found helpful. Thank You :)